I have another problem I need help with.
6+3y=y+4
Can you help? Thanks

Answer by checkley71(8403)

6+3y=y+4 or 3y-y=4-6 or 2y=-2 or y=-2/2 or y=-1
6+3(-1)=-1+4 or 6-3=-1+4 or 3=3